# Open-Set 3D object detection in LiDAR data as an Out-of-Distribution problem

### v1

@@ IV-B Two-Stage Methods and Input Choices
The two-stage methods use an MLP for OOD prediction. We first compare different inputs of the MLP using GT data (named ’Oracle MLP’ as in Section III-D):
@@
high-dimensional feature map (High-Dim): we can use higher-dimension (512) feature map representation from the base detector instead of the final lower-dimension (192) feature maps
@@
We show our experiments using the ground truth unknown objects for training in Table III. The results show that adding a 3x3 pooling layer and a focal loss was beneficial, compared to the work [revisiting], which only uses a binary cross-entropy loss, no pooling, high-dimensional feature maps, and feature interpolation. Despite a higher FPR-95, this is the only method that combines a higher precision on known objects and near-perfect precision on closed objects, which was outlined as a primary objective when seeing Open-Set as an OOD problem. We highlight this in the table, where using the input processing of [revisiting] leads to a known object precision of 70.5, and ours a 99.7 precision while maintaining an acceptable AUPR-E. The pooling smooths out the center interpolation errors from the detection model. For instance, in the Argoverse 2 dataset, a maxpooling filter enabled a feature aggregation operation with a radius of 2.1 meters. It substantially improved OOD performance by ensuring a more robust aggregation of surrounding point cloud data. While we trained our MLP models using the ground truth unknown objects in the previous section, we cannot access this information when training in real-world scenarios. Therefore, we need to generate pseudo-labels to train our model.
IV-C Methods for Generating Additional Unknown data for training the MLP
We then trained our MLP to detect OOD objects, using different generation methods of unknown objects, which we use as our ground truth.
IV-C1 Different Representations of Out-of-Distribution Objects
@@
We compare different representations of out-of-distribution objects for training, which we create through several means:
@@
Gaussian noise [du2022vos]: we perturb the extracted embedding of ID objects with Gaussian noise, adding a normal distribution to the extracted embeddings.
@@
External data: we incorporate ModelNet [modelnet], allowing us to introduce a diverse set of synthetic objects into our training regime, which we consider OOD. We insert between 15 and 25 objects, sampled randomly from the {’bench’, ’bookshelf’, ’chair’, ’cone’, ’desk’, ’flower pot’, ’lamp’, ’mantel’, ’night stand’, ’plant’, ’vase’, ’wardrobe’} class subset of ModelNet, and normalize their size to 1. We then sample 200 points at the surface of each object, followed by a grid sampling where we keep between 1 and 3 points with a grid size which we also randomize between 5 and 10, apply an intensity value to the points based on the intensity statistics of Argoverse 2, and finally resize the objects with a global scaling factor between 1 and 4.
@@
PointMixup [pointmixup]: we use this method to mix point clouds of different objects through point matching and interpolation to create new, hybrid representations. We expect objects created in such a fashion to contain local features present in both ID and OOD objects while creating novel appearances. Since [pointmixup] was designed for data augmentation, we chose different interpolation hyperparameters after testing variations. Each object with at least 5 points has a 20% probability of being replaced by a new version of that object generated as an interpolation between it and another object with at least 5 points, with interpolation factors randomly chosen in the range.
@@
Resizing: create out-of-distribution objects by resizing in-distribution (ID) known objects to modify their scales along all three axes, similarly to [revisiting].
@@
Top-K autolabelling approach [towardsopen]: the top-K most confident predictions from the base detector, which don’t overlap with ground truth ID objects, are used as OOD objects. In this approach, we keep the scans from the datasets containing real OOD distribution, as we expect these to become autolabelled. Since other methods do not rely on real OOD data for training the MLP, we remove those scans for other methods.
@@
IV-C2 Results
The results of our evaluation on the Argoverse 2 dataset, presented in Table IV, highlight the discrepancies in the MLP performance based on different generation methods of unknowns.
@@ IV-C2 Results
| Method | AUROC | FPR-95 | AUPR-E | AUPR-S |
| Gaussian Noise [du2022vos] | 66.5 | 84.5 | 3.7 | 98.9 |
@@
| PointMixUp†[pointmixup] | 72.1 | 72.4 | 7.4 | 99.1 |
| Resizing [revisiting] | 77.2 | 75.3 | 6.3 | 99.4 |
| Top-K Labelling (K=5) [towardsopen] | 81.9 | 68.2 | 8.8 | 99.6 |
| 2-Stage Oracle MLP | 86.8 | 51.1 | 20.0 | 99.7 |
ModelNet Injection, which incorporates synthetic objects from the ModelNet dataset as OOD data, showed better performance with an AUROC of 70.6 and an AUPR-E of 5.0. We believe the difficulties in OOD detection for this method stem from the synthetic nature of the dataset, which suggests that while other works suggest using insertions of instances as unknowns [UFO], one has to be careful about the quality of these instances and the domain differences between these and the real unknowns. PointMixUp struggled with a high FPR-95 of 77.6 and a low AUPR-E of 4.6, indicating challenges in detecting unknown objects despite the novel hybrid representations. The Top-K Labelling (K=5) method, which labels the top K most confident predictions not overlapping with ground truth ID objects as OOD, emerged as a notable performer. It achieved an AUROC of 81.9 and the highest AUPR-S of 99.6, demonstrating exceptional identification of known objects. Despite its higher FPR-95 of 68.2, Top-K Labelling showed promise in balancing the detection of known and unknown objects by leveraging real-world OOD distributions in the dataset. We experimented with different K values (K= ) and found that K=5, close to the average number of open class instances per frame, performed best across all metrics. Overall, the remaining gap between the two-stage oracle MLP and this highest-performing unknown generation process suggests that future works should focus on improving this pseudo-labelling process.
V CONCLUSIONS
Our work redefines the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution problem. Through this lens, OOD 3D object detection adds relevant information with regard to traditional object detection. In the process, we established a rigorous evaluation protocol and identified key evaluation parameters that we suggest highlighting in future works. Additionally, we show that two-stage methods show promise for OOD 3D Object Detection and underline the potential of autolabelling methods.

### v2

@@ III-E Ablation study on HD-OOD3D
The two-stage methods use an MLP for OOD prediction. We first compare different inputs of the MLP using the labels of unknown objets as previously:
@@
high-dimensional feature map (High-Dim): we use higher-dimension (512) feature map representation to extract the embeddings from the base detector instead of the final lower-dimension (192) feature maps
@@
We show our experiments using the ground truth unknown objects for training in Table III. The results show that adding a 3x3 pooling layer and a focal loss was beneficial, compared to the work [16], which only uses a binary cross-entropy loss, no pooling, high-dimensional feature maps, and feature interpolation. Despite a higher FPR-95, this is the only method that combines a higher precision on known objects and near-perfect precision on closed objects. For instance, in the Argoverse 2 dataset, a maxpooling filter enabled a feature aggregation operation with a radius of 2.1 meters. It substantially improved OOD performance by ensuring a more robust aggregation of surrounding point cloud data.
IV Analysis of the Evaluation Protocol for 3D OOD
Here, we examine how external factors and hyperparameter choices influence OOD detection performance because these elements directly affect both the quality and reliability of the detection metrics. For instance, the characteristics of the dataset—such as point cloud density, class diversity, and the presence of noise—can alter how many true and false positives are identified. Similarly, hyperparameters like the distance threshold for matching predictions, the detection score cutoff, and the criteria used for sorting predictions determine which detections are considered valid. Even small adjustments in these parameters can lead to significant shifts in performance metrics, affecting the reproducibility and fairness of the evaluation. Therefore, we examine and compare these choices.
IV-A Metrics and Matching
As previously described, we use the AUROC, FPR-95, AUPR-E (with unknown classes as positives) and AUPR-S (with known classes as positives) to assess the model’s ability to distinguish between known and unknown objects. We also describe our matching algorithm, detailed in Algorithm 1, which is crucial for computing True Positives (TP), False Positives (FP), True Negatives (TN), and False Negatives (FN). This algorithm, which includes essential hyperparameters, plays a significant role in determining the final evaluation metrics, mainly by ignoring any predicted box whose center is not within a fixed range of any non-assigned ground truth.
We use the ground truth data to train the MLP with the actual annotated ’unknown’ classes, like in Section III. We demonstrate that many external factors could impact the performance and resulting metrics, which are not disclosed in [16]. We identify several key parameters that significantly influence the effectiveness of HD-OOD3D when applied to GT data, allowing us to examine the influence of these parameters independently of the actual performance of the MLP.
For the two-stage family of methods, we train an MLP after freezing the initial detector Centerpoint, like in the previous section, and use the same MLP architecture. For ease of reading, we include the relevant experiments from our two-stage HD-OOD3D in the following.
IV-B Hyperparameter Analysis
IV-B1 Evaluation set
Using the ”Open” Scans subset, which includes only scans with unknown objects, significantly improved the AUPR-E metric from 12.6 to 20.0 for HD-OOD3D. Due to the known/unknown class imbalance, using all scans adds many more potential false positives without any unknown ground truth, hindering all methods across the board. Consequently, our final evaluation used the ”open” subset—only the evaluation scans containing at least one unknown object—to better assess performance. It is important to note that, as shown in Table IV, the chosen hyperparameters do not maximize the performance of HD-OOD3D in isolation, but rather provide a balanced and fair benchmark for all methods.
IV-B2 Distance and Threshold Settings for OOD Detection
A lower threshold in the detection module (e.g., 0.3) allows more predicted objects to pass through, reducing misses, especially for known objects. However, this comes at the cost of an increased likelihood of false positives among the unknown detections. A trade-off exists: while a lower threshold increases the detection rate of unknown objects, it also results in more potential false positives. Therefore, a threshold of 0.3, paired with a 2.0m distance for box matching, strikes a balance between minimizing misses for known objects and enhancing the detection of unknown objects, as evidenced by the metrics in Table IV.
| Hits (U)% | Hits (C)% | AUROC | FPR-95 | AUPR-E | AUPR-S | ||
| 0.05 | 0.5m | 22.8 | 49.3 | 91.0 | 43.7 | 22.2 | 99.8 |
| 2.0m | 38.9 | 55.4 | 84.9 | 62.8 | 17.5 | 99.5 | |
| 0.3 | 0.5m | 14.4 | 45.3 | 92.3 | 37.4 | 23.2 | 99.9 |
| 2.0m | 26.2 | 50.6 | 88.6 | 51.1 | 20.0 | 99.7 | |
| 0.5 | 0.5m | 7.6 | 39.0 | 93.6 | 32.6 | 24.8 | 99.9 |
| 2.0m | 11.7 | 42.4 | 91.4 | 38.7 | 22.1 | 99.9 |
IV-B3 Sorting method
From Table IV, we notice widely different results from the traditional OOD metrics for any fixed method, depending on the evaluation hyperparameters. The choice of detection score threshold and distance threshold is crucial for balancing the precision and recall of OOD detection. Both and determine how detections are matched to ther ground truth: larger threshold includes more detections but risk misassigning predictions. Rather than focusing on optimizing the performance of our method, our primary goal was to refine the evaluation protocol to ensure a fair and robust comparison. Through systematic evaluation, we found that setting and m provides the best trade-off, maximizing recall for OOD objects while minimizing degradation on known classes. These choices allow for a more reliable assessment of OOD detection performance across different methods. We suggest that future works consider the recall of the base detector and explicitly specify these crucial evaluation hyperparameters to facilitate reproducibility and fair benchmarking.
V Methods for Generating Additional Unknown data for training the MLP
In the previous sections, our MLP models were trained using ground truth labels for unknown objects. However, in real-world scenarios, manually annotating every unknown object is infeasible due to the diversity and unpredictability of OOD instances. Unlike in controlled datasets like Argoverse 2, where unknown objects can be labeled explicitly, a practical deployment scenario requires detecting OOD objects without prior supervision. To address this challenge, we propose generating pseudo-labels for unknown objects, enabling training without explicit ground truth annotations.
V-A Different Representations of Out-of-Distribution Objects
@@
In contrast to the single-generation approach used in MMOOD3D [16], we propose several alternative methods to synthesize out-of-distribution (OOD) objects, specifically a Top-K labelling approach. We compare these different representations of out-of-distribution objects for training, which we create through several means:
@@
Gaussian noise [9]: we perturb the extracted embedding of ID objects with Gaussian noise, adding a normal distribution to the extracted embeddings.
@@
External data: we incorporate ModelNet [22], allowing us to introduce a diverse set of synthetic objects into our training regime, which we consider OOD. We insert between 15 and 25 objects, sampled randomly from the {’bench’, ’bookshelf’, ’chair’, ’cone’, ’desk’, ’flower pot’, ’lamp’, ’mantel’, ’night stand’, ’plant’, ’vase’, ’wardrobe’} class subset of ModelNet, and normalize their size to 1. We then sample 200 points at the surface of each object, followed by a grid sampling where we keep between 1 and 3 points with a grid size which we also randomize between 5 and 10, apply an intensity value to the points based on the intensity statistics of Argoverse 2, and finally resize the objects with a global scaling factor between 1 and 4.
@@
PointMixup [23]: we use this method to mix point clouds of different objects through point matching and interpolation to create new, hybrid representations. We expect objects created in such a fashion to contain local features present in both ID and OOD objects while creating novel appearances. Since [23] was designed for data augmentation, we chose different interpolation hyperparameters after testing variations. Each object with at least 5 points has a 20% probability of being replaced by a new version of that object generated as an interpolation between it and another object with at least 5 points, with interpolation factors randomly chosen in the range.
@@
Resizing: create out-of-distribution objects by resizing in-distribution (ID) known objects to modify their scales along all three axes, similarly to [16]. Each known object with at least 5 points has a 50% probability of being resized, where for each axis, we randomly apply a scaling factor between and .
@@
Our 3D Top-K autolabelling : inpired by [3] for open 2d objet detection in images, we propose in 3D to use the top-K most confident predictions from the base detector, which don’t overlap with ground truth known objects, as OOD objects. In this approach, we keep the scans from the datasets containing real OOD distribution, as we expect these to become autolabelled. Since other methods do not rely on real OOD data for training the MLP, we remove those LiDAR scans for other methods.
@@
V-B Results
The results of our evaluation on the Argoverse 2 dataset, presented in Table V, highlight the discrepancies in the MLP performance based on different generation methods of unknowns.
@@ V-B Results
| Unsupervised Unknowns | AUROC | FPR-95 | AUPR-E | AUPR-S |
| Gaussian Noise [9] | 66.5 | 84.5 | 3.7 | 98.9 |
@@
| PointMixUp†[23] | 72.1 | 72.4 | 7.4 | 99.1 |
| Resizing from MMOOD3D[16] | 77.2 | 75.3 | 6.3 | 99.4 |
| Top-5 Labelling (Ours) | 81.9 | 68.2 | 8.8 | 99.6 |
| GT Unknowns | 86.8 | 51.1 | 20.0 | 99.7 |
ModelNet Injection, which incorporates synthetic objects from the ModelNet dataset as OOD data, showed better performance with an AUROC of 70.6 and an AUPR-E of 5.0. We believe the difficulties in OOD detection for this method stem from the synthetic nature of the dataset, which suggests that while this method might be interesting [14], one has to be careful about the quality of these instances and the domain differences between these and the real unknowns. PointMixUp struggled with a high FPR-95 of 77.6 and a low AUPR-E of 4.6, indicating challenges in detecting unknown objects despite the novel hybrid representations. The Top-K Labelling (K=5) method, which labels the top 5 most confident predictions not overlapping with ground truth known objects as OOD, emerged as a notable performer. It achieved an AUROC of 81.9 and the highest AUPR-S of 99.6, demonstrating good identification of known objects. Despite its higher FPR-95 of 68.2, Top-K Labelling showed promise in balancing the detection of known and unknown objects by leveraging real-world OOD distributions in the dataset. We experimented with different K values (K= ) and found that K=5, close to the average number of unknown class instances per frame, performed best across all metrics. Overall, the remaining gap between the two-stage oracle MLP and this highest-performing unknown generation method suggests that future works should focus on improving this pseudo-labelling process.
VI CONCLUSIONS
In this work, we presented a new two-stage method for detecting unknown objects called HD-OOD3D. By exploiting a sufficiently large dataset such as Argoverse 2 to train the unknowns in a supervised manner, we demonstrated the superiority of two-stage methods over single-stage methods, particularly our proposed HD-OOD3D, which utilizes more input representations for better object categorization. We then broke down the classic evaluation protocol for unknown objects to highlight the importance of certain hyperparameter choices for a fair assessment of methods. Finally, scaling the learning of unknown objects can only be achieved through unsupervised training. We compared different approaches for generating unknowns by using only the annotations of known objects from Argoverse 2. We showed that top-5 auto-labelling offers more promising performance perspectives compared to resizing. Ultimately, the task of detecting unknown objects remains a challenging one, with current performance far from what would enable deployment on robotics platforms, making it a field that still requires further research to reach maturity.